OpenStudy (anonymous):

What is 10^-3 = 1/1000 in logarithmic form?

OpenStudy (e.cociuba):

Isn't the prob supposed to look like this 1/(10)^-3 = 1000??

OpenStudy (anonymous):

no @e.cociuba it looks like this

OpenStudy (anonymous):

@LazyBoy

OpenStudy (anonymous):

\[BASE^{EXPONENT} = ANSWER\] \[\log_{BASE}ANSWER = EXPONENT\]

pooja195 (pooja195):

in logarithmic form log 1 - (-3log10) = log 1000 that is log 1 + 3log10 = log 1000 (log 1 base 10 = 0 + 3 log 10 base 10 = 3 * 1 = 3 so 0 + 3 = 3 and log 1000 base 10 = 3 , All in base 10) ie: log (1/10) base 1000= -3